If 3 times the square of a number is added to the number itself, the result is 14. Find the number? I need help on how to set up the problem.
Thank you

Answer by scott8148(6628)

n + 3n^2 = 14

3n^2 + n - 14 = 0

factoring ___ (3n + 7)(n - 2) = 0